Determine the mass of oxygen required to completely burn 10.0 g of propane (C3H8).
1
Expert's answer
2020-05-23T12:24:14-0400

The balanced chemical equation for the combustion of propane is: 

C3H8(g)+5O2(g)→3CO2(g)+4H2O(g)

given amount of propane is 10 grams

molecular weight of propane=44 gram/mol


so no. of moles of propane = 10/44 moles


since 1 moles of propane reacts with 5 moles of oxygen.

so no. of moles of oxygen required = 5* no of moles of propane

no of moles of oxygen required= 5*10/44 moles


mass of oxygen required= no. of moles * Molecular weight of oxygen

mass of oxygen required= 5*10/44*32 grams

=36.36 grams